OpenStudy (anonymous):

integral sec^4 (x/2) dx from 0 to pi/2

OpenStudy (anonymous):

you deleted the older post..

OpenStudy (anonymous):

\[\int\limits_{?}^{?}\sec^4(\frac{x}{2})dx\]

OpenStudy (anonymous):

sec^2 * sec^2(x) and change one into tangent

OpenStudy (anonymous):

or tan i mean

OpenStudy (anonymous):

then you will have a derivative and it's function

OpenStudy (anonymous):

anyway apply integration by parts and simplify\[I = 2\sec^2(x/2) \tan(x/2) - 4I + 4\int\limits\limits_{}^{} \tan(x/2) \sec^2(2x/2)dx\] to get

OpenStudy (anonymous):

\[\cos^2+\sin^2=1\] \[1+\tan^2=\sec^2\]

OpenStudy (anonymous):

eww that is ugly...

OpenStudy (anonymous):

\[5I = 2\sec^2(x/2) \tan(x/2)+ 8\sec^2(x/2)\] on applying limits we get -3/5

OpenStudy (anonymous):

sec^4(x/2) dx u=x/2 du=1/2 dx 2du=dx 2/cos^4 (u) du

OpenStudy (anonymous):

double susdtitution

OpenStudy (anonymous):

you could do double substitution but i still believe splitting the sec up and then creating two integrals from that

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

like this\[\int\limits_{?}^{?}\tan^2(x/2)\sec^2(x)dx +\int\limits_{?}^{?}\sec^2(x)dx\]

OpenStudy (anonymous):

i mean x/2 for the other one

OpenStudy (anonymous):

u = tan(x/2) du = 1/2(sec^2(x/2)dx

OpenStudy (anonymous):

2du=sec^2(x/2)dx and then you have just integral of u^2

OpenStudy (anonymous):

the second one you can just take the integral it will just be tan(x/2)

OpenStudy (anonymous):

or double substitution.,..... but parts seems a little insane. it'd work but that would be a lot of evaluating